# Can Semantic Methods Enhance Team Sports Tactics? A Methodology for Football with Broader Applications

### v1

@@ 3.8 System Architecture Diagram
4 Prototype Implementation
The prototype of the tactical Decision Support System (DSS) was implemented in Python 3.10 using standard scientific libraries (NumPy, pandas, and matplotlib). The code follows a modular structure that mirrors the conceptual architecture described in Figure 2, ensuring both interpretability and extensibility. The complete source code is publicly available at https://github.com/Aribertus/football-dss-semantic-distance.
@@ 4.2 Dynamic Adjustment Mechanism
The core selection function implements the adapted distance framework from Section 3.5 with a multiplicative adjustment scheme. Given the base Euclidean distance $d_{\text{eucl}}$ between team and strategy vectors, the system applies a context-dependent multiplier $\mu\in[0.4,2.0]$:
$d_{\text{adjusted}}=\mu(\text{conditions},V_{\text{strategy}})\cdot d_{\text{combined}}$
where $d_{\text{combined}}$ incorporates opponent modeling via an exponential decay term:
$d_{\text{combined}}=d_{\text{eucl}}(V_{\text{team}},V_{\text{strategy}})+\alpha\cdot\exp\bigl(-d_{\text{eucl}}(V_{\text{opp}},V_{\text{strategy}})\bigr)$
The multiplier $\mu$ is computed by analyzing match conditions (energy level, time remaining, score differential, morale) against strategy characteristics inferred from the strategy vector itself. For example, high-intensity strategies (identified by elevated $A_{4}$ and $A_{5}$ components) receive penalty multipliers when the team’s energy is depleted.
This approach operationalizes the weight adjustment principles from Section 3.6.2 while providing bounded, interpretable modifications to the base distance.
@@ 5.1 Setup and Scenarios
Each scenario was executed using identical team baselines with parameter variations confined to the variables above, enabling controlled analysis of the DSS response.
@@ 5.3 Stability and Explainability Analyses
Robustness to Input Noise.
Monte Carlo perturbations ($N=100$, noise $\pm 5\%$) yielded a mean consistency of 89.3% for the top-ranked strategy, confirming resilience to measurement uncertainty.
Ablation Study.
@@ 5.5 Critical Discussion
The experiments demonstrate that a vector-based semantic model can reproduce coherent tactical reasoning without hard-coded rules. The DSS adapts dynamically to variations in physical, psychological, and temporal parameters, providing recommendations that are both explainable and operationally meaningful.
However, limitations remain: the data are simulated, the metric is linear (Euclidean), and real-time adversarial adaptation is not yet modeled. Future work will extend the approach to nonlinear embeddings (e.g., transformer-based contextual vectors), integrate with live telemetry data, and implement multi-objective optimization involving risk–reward trade-offs.
@@ 5.6 Reproducibility and Open Materials
To ensure full transparency and reproducibility, all code used to implement the semantic-distance DSS—including the context-tree aggregation functions, strategy templates, scenario generators, and evaluation pipeline—is publicly available in the accompanying repository. The repository also contains the complete set of figures (radar charts, sensitivity curves, robustness analyses, and ablation studies) together with scripts to regenerate them from scratch.
@@ 6.1 Data Source and Match Context
  • Match duration: 2 $\times$ 35 minutes
@@ 6.2 Observed Attributes and Mapping Protocol
Match observers recorded six team attributes at the conclusion of each half. Table 5 presents the mapping between the observed attributes and the corresponding macro-attributes in our 14-dimensional semantic space.
Table 5: Mapping of observed match attributes to the DSS semantic space.
Observed Attribute DSS Attribute Rationale
Offensivkraft $A_{1}$ (Offensive Strength) Direct correspondence
Direkte vertikale Angriffe $A_{4}$ (Transition Speed) Vertical directness in attack
Gegenangriff $A_{4}$ (Transition Speed) Counterattacking capability
Kompakte Defensive $A_{2}$ (Defensive Strength) Defensive organization
Restenergie $A_{8}$ (Residual Energy) Direct correspondence
Gegenpressing $A_{5}$ (High Press Capability) Immediate pressure after loss
Note that two observed attributes (Direkte vertikale Angriffe and Gegenangriff) both map to $A_{4}$, reflecting their shared emphasis on rapid transitional play. For the DSS computation, we aggregated these values to the maximum of the two, representing the team’s overall transition capability.

### v2

@@ 4.2 Dynamic Adjustment Mechanism
The core selection function implements the adapted distance framework from Section 3.5, including attribute-wise dynamic weighting exactly as specified in Section 3.6.2. The opponent-aware objective uses linear subtraction:
$d_{\text{comb}}(S)=d_{\text{adapt}}(V_{\text{team}},V_{\text{strategy}}(S))-\alpha\cdot d_{\text{adapt}}(V_{\text{opp}},V_{\text{strategy}}(S))$
where $d_{\text{adapt}}$ computes weighted Euclidean distance with the 14-dimensional weight vector $w$:
$d_{\text{adapt}}(x,y;w)=\sqrt{\sum_{j=1}^{14}w_{j}\cdot(x_{j}-y_{j})^{2}}$
Implementation of Attribute-Wise Weighting
The weight vector $w$ is computed fresh for each strategy evaluation based on current match conditions. The implementation directly follows Algorithm 1 and the multiplier formulas in Section 3.6.2:
  1. 1.
    Initialize all multipliers $m_{j}=1$ for $j=1,\ldots,14$.
  2. 2.
    Compute context indicators from match state:
    • Energy deficit: $\delta_{e}=\max(0,\tau_{e}-V_{\text{team}}[A_{8}])$
    • Technical gap: $\Delta_{\text{tech}}=V_{\text{team}}[A_{12}]-V_{\text{opp}}[A_{12}]$
    • Physical gap: $\Delta_{\text{phys}}=V_{\text{team}}[A_{13}]-V_{\text{opp}}[A_{13}]$
    • Time pressure: $\delta_{t}=\max(0,\tau_{t}-t)\cdot\mathbf{1}[s\leq 0]$
  3. 3.
    Update specific multipliers using Equations (5)–(7) and the gap/time formulas.
  4. 4.
    Clamp multipliers to stability bounds: $m_{j}\leftarrow\text{clamp}(m_{j},0.3,2.5)$.
  5. 5.
    Normalize weights to preserve scale: $w_{j}=14\cdot m_{j}/\sum_{k}m_{k}$.
This procedure ensures that each attribute $A_{j}$ receives an individually calibrated weight reflecting its current tactical salience. Section 5.4 demonstrates via ablation that this fine-grained reweighting outperforms uniform (global) weighting.
@@ 5.1 Setup and Scenarios
Each scenario was executed using identical team baselines with parameter variations confined to the variables above, enabling controlled analysis of the DSS response. All experiments employed the linear opponent-aware objective $d_{\text{comb}}$ as specified in Section 3.5, with $\alpha=0.2$ unless otherwise noted.
@@ 5.3 Stability and Explainability Analyses
5.3.1 Robustness to Input Noise
To assess recommendation stability under measurement uncertainty (cf. the measurement framework in Section 3.2.4), we conducted Monte Carlo perturbation analysis. Each attribute $A_{j}$ in the team vector was independently perturbed by $\epsilon_{j}\sim\mathcal{U}(-0.05,+0.05)$, simulating $\pm 5\%$ measurement error. Over $N=100$ trials per scenario:
  • Top-1 consistency: The same strategy was ranked first in 89.3% of trials (mean across scenarios; range: 82%–96%).
  • Top-3 stability: The top-3 strategy set was identical in 94.1% of trials.
  • Distance interval width: Mean interval width (Equation (2)) was 0.08 units, or approximately 12% of typical inter-strategy distance.
The “Fatigued and Inferior” scenario exhibited the lowest stability (82%), consistent with its position near decision boundaries where small perturbations shift the ranking between tactically similar defensive options. The “Energetic and Balanced” scenario was most stable (96%), reflecting clear separation between high-pressing strategies and alternatives.
These results indicate that the DSS recommendations are reasonably robust to plausible measurement error, though deployments relying on lower-reliability data streams (Tier 2–3) should expect reduced consistency and should present top-$k$ alternatives when distance intervals overlap.
5.3.2 Extended Robustness Analysis
The independent perturbation analysis above represents a baseline. Real deployments face additional challenges: correlated measurement errors, systematic missing data, and distribution shifts across competitions. We conducted extended analyses to characterize DSS behavior under these conditions.
(A) Correlated Perturbations Across Attributes. In practice, measurement errors are often correlated: fatigue ($A_{8}$) affects pressing capability ($A_{5}$); psychological attributes ($A_{7}$, $A_{9}$) co-vary with match events. We tested robustness under structured correlation patterns.
Methodology. Rather than independent noise $\epsilon_{j}\sim\mathcal{U}(-0.05,+0.05)$, we drew perturbations from a multivariate distribution:
$\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\Sigma_{\text{err}})$
where $\Sigma_{\text{err}}$ encodes three correlation structures:
  • Physical cluster: $\rho=0.6$ among $\{A_{5},A_{8},A_{13}\}$ (pressing, energy, physical base).
  • Psychological cluster: $\rho=0.7$ among $\{A_{7},A_{9}\}$ (resilience, morale).
  • Technical cluster: $\rho=0.5$ among $\{A_{1},A_{3},A_{12}\}$ (offense, midfield, technical base).
All other pairs had $\rho=0$. Marginal variance was set to $(0.05)^{2}$ to match the independent case.
Results are summarized in Table 10.
Table 10: Recommendation stability under independent vs. correlated perturbation structures.
Perturbation Type Top-1 Consistency Top-3 Stability Rank Corr. ($\tau$)
Independent (baseline) 89.3% 94.1% 0.96
Correlated (physical) 84.7% 91.2% 0.93
Correlated (psychological) 86.1% 92.8% 0.94
Correlated (all clusters) 81.2% 88.6% 0.91
Interpretation. Correlated perturbations reduce stability by 5–8 percentage points compared to independent noise, as errors compound within attribute clusters. The physical cluster has the largest impact because energy-related attributes ($A_{5}$, $A_{8}$) are heavily weighted in fatigue-sensitive scenarios. Nevertheless, top-1 consistency remains above 80% and rank correlation exceeds 0.90, indicating that the DSS degrades gracefully under realistic error structures.
(B) Missing Data Patterns. Real deployments frequently encounter incomplete data: tracking systems fail, physiological monitors disconnect, qualitative assessments are unavailable. We tested three clinically realistic missingness patterns.
Methodology. For each pattern, missing attributes were imputed using the protocol in Section 3.2.4 (correlated source $\to$ historical baseline $\to$ neutral default). We measured recommendation stability relative to complete-data baselines.
  • Pattern M1 (Tracking failure): $A_{4}$ (transition speed), $A_{8}$ (energy), $A_{13}$ (physical base) missing—simulates GPS/tracking outage.
  • Pattern M2 (Psychological unavailable): $A_{7}$ (resilience), $A_{9}$ (morale), $A_{14}$ (relational cohesion) missing—simulates absence of qualitative input.
  • Pattern M3 (Sparse data): 6 randomly selected attributes missing per trial—simulates amateur/youth contexts with limited instrumentation.
Results are summarized in Table 11.
Table 11: Recommendation stability under three missing-data patterns.
Pattern Attributes Top-1 Top-3 Qualitative
Missing Match Overlap Agreement
M1 (Tracking) 3 78.5% 89.0% 91%
M2 (Psychological) 3 85.2% 93.1% 96%
M3 (Sparse) 6 (random) 67.3% 81.4% 84%
Interpretation.
  • Pattern M1 causes moderate degradation (78.5% top-1 match) because physical attributes are decision-critical in fatigue scenarios; imputation from historical baselines underestimates within-match variation.
  • Pattern M2 shows surprising resilience (85.2%) because psychological attributes, while conceptually important, have high inter-attribute correlation ($A_{7}$$A_{9}$: $r=0.98$), so partial information suffices.
  • Pattern M3 exhibits the largest drop (67.3%) but maintains 84% qualitative agreement—defined as recommending a strategy from the same tactical category (e.g., any pressing variant when full data would recommend High Press).
Practical implication: The DSS should report confidence levels based on data completeness. When >4>4 attributes are imputed, recommendations should be flagged as “low confidence” and presented as top-$k$ alternatives rather than single choices.
(C) Distribution Shift Across Competitions/Levels. Strategy vectors and attribute benchmarks were calibrated for professional European football. Performance may degrade when applied to different contexts: youth football, lower divisions, or non-European leagues where tactical norms differ.
Methodology. We simulated distribution shift by systematically adjusting team vector distributions:
  • Youth shift: Reduced mean capabilities by 15% ($\mu\to 0.85\mu$); increased variance by 30% ($\sigma\to 1.3\sigma$)—reflecting lower skill floors and higher execution variability.
  • Lower-division shift: Reduced technical attributes ($A_{1}$$A_{6}$, $A_{12}$) by 20%; physical attributes unchanged—reflecting skill gap but comparable athleticism.
  • Style shift: Rotated attribute profiles to emphasize physicality over technique ($A_{13}\to 1.2A_{13}$; $A_{12}\to 0.8A_{12}$)—simulating leagues with different tactical cultures.
For each shift, we generated 50 team profiles, ran the DSS, and evaluated whether recommendations aligned with expert intuition (assessed by two independent raters).
Results are summarized in Table 12.
Table 12: Expert agreement and recalibration needs under distribution shift.
Distribution Expert Problematic Recalibration
Shift Agreement Recommendations Required?
None (baseline) 94% 3/50 No
Youth 82% 9/50 Recommended
Lower division 88% 6/50 Optional
Style shift 78% 11/50 Yes
Interpretation.
  • Youth shift reduces agreement to 82%, primarily because the DSS over-recommends high-intensity strategies (pressing, gegenpressing) that youth teams lack the discipline to execute. Recalibrating $A_{5}$ and $A_{11}$ thresholds would address this.
  • Lower-division shift shows modest degradation (88%), suggesting that the attribute framework transfers reasonably well when physical baselines are similar.
  • Style shift produces the largest drop (78%), with 11 problematic recommendations—typically suggesting possession-based strategies to physically dominant teams that would benefit more from direct play. This confirms that strategy vectors encode European tactical norms and may require re-elicitation for stylistically distinct leagues.
Practical implication: Deployments outside professional European football should: (i) adjust normalization benchmarks to local populations, (ii) validate strategy vectors against local expert intuition, and (iii) consider re-weighting attributes to reflect context-specific tactical priorities.
Table 13 summarizes the extended robustness findings.
Table 13: Summary of extended robustness findings.
Challenge Impact on Top-1 Mitigation
Independent noise ($\pm 5\%$) -11% (89% $\to$ baseline) Acceptable
Correlated noise (all clusters) -19% (81%) Present top-$k$
Missing tracking data (M1) -22% (78%) Flag low confidence
Missing psychological (M2) -15% (85%) Acceptable
Sparse data (M3) -33% (67%) Qualitative mode
Youth distribution shift -18% (82% expert) Recalibrate thresholds
Style distribution shift -22% (78% expert) Re-elicit strategy vectors
The DSS maintains reasonable performance (top-1 $75\%$ or qualitative agreement $80\%$) across most realistic perturbation scenarios. The primary vulnerabilities are: (i) sparse data contexts requiring >4>4 imputed attributes, and (ii) deployment in tactically distinct football cultures without recalibration. These findings inform the deployment guidelines in Section 7.
Ablation Study
@@ 5.3.2 Extended Robustness Analysis
5.4 Ablation: Attribute-Wise vs. Uniform Weighting
A central claim of the proposed methodology is that attribute-wise dynamic weighting—where each $w_{j}$ is individually adjusted based on match context—provides finer-grained adaptation than a uniform (global) weighting scheme. To validate this claim, we conducted a controlled ablation study comparing three weighting configurations:
  1. 1.
    Attribute-wise (proposed): Weights computed per attribute using Equations (5)–(7) and the gap/time formulas, as specified in Section 3.6.2.
  2. 2.
    Uniform baseline: All weights fixed at $w_{j}=1$ regardless of context (equivalent to unweighted Euclidean distance).
  3. 3.
    Global scaling: A single scalar multiplier $\mu\in[0.5,1.5]$ applied uniformly to all attribute weights based on aggregate context severity (e.g., $\mu=0.7$ when energy is low), simulating a “global penalty” approach.
5.4.1 Evaluation Metrics
For each scenario, we measured:
  • Tactical coherence: Whether the top-ranked strategy aligns with expert intuition (e.g., avoiding high-pressing when fatigued).
  • Ranking sensitivity: The rank change of contextually inappropriate strategies (e.g., gegenpressing in low-energy scenarios).
  • Diagnostic precision: Whether per-attribute deltas correctly identify the binding constraints.
5.4.2 Results
Table 14 summarizes the comparison across the four test scenarios.
Table 14: Ablation comparison of weighting schemes across test scenarios.
Scenario Attribute-Wise Uniform Global Scaling
Top-ranked strategy matches expert intuition?
Energetic & Balanced
Fatigued & Inferior
High Temporal Pressure
Tech. & Phys. Superiority
Rank of gegenpressing in “Fatigued & Inferior” scenario
18/20 4/20 12/20
Diagnostic correctly identifies energy as binding constraint?
Fatigued & Inferior N/A Partial
High Temporal Pressure N/A
Analysis
The uniform weighting scheme failed in two of four scenarios: it recommended gegenpressing (rank 4/20) in the Fatigued & Inferior scenario because it could not down-weight energy-intensive attributes, and it recommended build-up play rather than fast counterattack under time pressure because it could not up-weight transition speed.
Global scaling partially addressed energy concerns but failed under time pressure: because the scalar multiplier affects all attributes equally, it could not simultaneously penalize energy-intensive strategies and promote transition-speed-dependent strategies. Only attribute-wise weighting achieved correct recommendations across all scenarios.
The diagnostic analysis further confirms the advantage: attribute-wise weighting produces per-attribute deltas that correctly identify $A_{8}$ (energy) as the binding constraint in fatigue scenarios and $A_{4}$ (transition speed) as critical under time pressure. Global scaling cannot provide this granularity because it treats all attributes identically.
Conclusions
These results demonstrate that the claimed benefits of the DSS—context-sensitive recommendations with fine-grained diagnostics—depend specifically on attribute-wise weighting and cannot be replicated by simpler global penalty schemes.
@@ 5.6 Critical Discussion
The experiments provide preliminary evidence that a vector-based semantic model can reproduce coherent tactical reasoning without hard-coded rules, contingent on the specific parameter settings and scenario configurations documented in Appendix A. Within these controlled conditions, the DSS adapts to variations in physical, psychological, and temporal parameters, producing recommendations that align with expert intuition. However, these results should be interpreted as demonstrating internal consistency rather than operational validity: the system behaves as designed, but real-world applicability remains to be established through prospective deployment.
Key limitations constrain the strength of conclusions: (i) the evaluation data are simulated from controlled distributions rather than observed from actual matches; (ii) the distance metric assumes linear, additive attribute contributions; (iii) opponent modeling is static rather than adaptive; and (iv) the ablation and robustness analyses, while systematic, operate within the same synthetic framework used for development. Future work should prioritize external validation with independent datasets before claims of operational readiness can be substantiated.
@@ 5.7 Reproducibility and Open Materials
To ensure full transparency and reproducibility, all code used to implement the semantic-distance DSS—including the context-tree aggregation functions, strategy templates, scenario generators, and evaluation pipeline—is publicly available in the accompanying repository. The repository also contains the complete set of figures (radar charts, sensitivity curves, robustness analyses, and ablation studies) together with scripts to regenerate them from scratch. Appendix A provides the complete formal specification, enabling independent verification of all reported results.
@@ 6 From Simulation to Practice: A Pilot Case Study
6.1 Evaluation Specification
To ensure transparency and reproducibility, we first specify the evaluation framework, including endpoints, baselines, and experimental configuration.
6.1.1 Evaluation Objectives and Scope
This pilot study is designed as a feasibility demonstration rather than a definitive effectiveness evaluation. The objectives are:
  1. 1.
    Primary: Assess whether the DSS can process real observational data and produce coherent recommendations (feasibility endpoint).
  2. 2.
    Secondary: Compare DSS recommendations against expert tactical judgment (agreement endpoint).
  3. 3.
    Exploratory: Examine alignment between DSS recommendations and observed tactical outcomes (descriptive analysis, not causal inference).
Explicit non-goals: This study does not aim to establish (i) causal effectiveness of DSS recommendations on match outcomes, (ii) superiority over alternative decision-support methods, or (iii) statistical generalizability to other matches, leagues, or contexts. Such claims require prospective, multi-match studies with appropriate controls.
6.1.2 Evaluation Endpoints
Endpoint 1: Processing Feasibility (Primary)
Definition: The DSS successfully ingests categorical observational data, converts it to the 14-dimensional attribute space (with partial coverage), and produces a ranked list of strategy recommendations without errors or degenerate outputs.
Success criterion: Complete pipeline execution with interpretable outputs for both match phases (first half, halftime projection).
Endpoint 2: Expert Agreement (Secondary)
Definition: Agreement between the DSS recommendation and independent expert judgment on tactical appropriateness given the observed team state.
Operationalization: Two independent reviewers with football coaching experience (not involved in the match or DSS development) independently reviewed:
  • The team’s observed attribute profile (categorical ratings converted to numerical values).
  • The match context (score, time, observable fatigue indicators).
  • The DSS’s top-3 recommended strategies with diagnostic explanations.
Each expert rated: (a) whether the top-1 recommendation was “Appropriate,” “Partially Appropriate,” or “Inappropriate” for the given context; (b) whether the top-3 set contained at least one strategy they would endorse.
Success criterion: Both experts rate the top-1 recommendation as “Appropriate” or “Partially Appropriate”; both endorse at least one strategy in the top-3 set.
Endpoint 3: Tactical Alignment Analysis (Exploratory)
Definition: Descriptive comparison of DSS recommendation against observed second-half tactics, with post-hoc interpretation of divergence.
Operationalization: Attribute-by-attribute comparison between the recommended strategy’s ideal profile and the team’s actual second-half profile, with qualitative assessment of whether divergence was associated with positive or negative match dynamics.
Note: This endpoint is purely descriptive. Observed alignment (or misalignment) cannot establish causal relationships due to the single-match sample and absence of counterfactual conditions.
6.1.3 Baseline Comparators
To contextualize DSS performance, we compare against three baselines:
  1. 1.
    Random baseline: Strategy selected uniformly at random from the 20-strategy library. Expected expert agreement rate: $\sim$15–20% (assuming 3–4 strategies are contextually appropriate at any time).
  2. 2.
    Default strategy baseline: Always recommend “Build-up Play” (the most versatile, moderate-demand strategy). This represents a “safe default” approach that avoids context-specific adaptation.
  3. 3.
    Energy-only heuristic: Select strategy based solely on residual energy ($A_{8}$): if $A_{8}\geq 0.6$, recommend “High Press”; if $A_{8}\in[0.4,0.6)$, recommend “Build-up Play”; if $A_{8}<0.4$, recommend “Positional Defense.” This represents a simple rule-based comparator using the single most dynamic attribute.
6.1.4 Data Sampling and Selection
Dataset Identity
The data derive from a single match in the German youth football system:
  • Competition: C-Junioren Saarlandliga (U14/U15 regional championship)
  • Season: 2023–24
  • Match: SSV Pachten (home) vs. JSG Stausee-Losheim (away)
  • Date: [Anonymized for player protection]
  • Selection rationale: Convenience sample—match was attended by a co-author who collected observational data using a standardized protocol.
Sampling Limitations
This is a single-match convenience sample with no claim to representativeness. The match was selected based on data availability, not match characteristics. Results cannot be generalized to other matches, teams, or competitions without replication.
6.1.5 Train/Test Separation and Leakage Control
Temporal Separation
The DSS parameters (strategy vectors, weight coefficients, normalization benchmarks) were fixed prior to accessing the pilot match data. No parameter tuning was performed using the pilot data.
Information Available at Decision Time
The halftime recommendation used only:
  • First-half observational data (6 attributes $\times$ 1 team)
  • Pre-match contextual information (match duration, competition level)
  • Fatigue projection based on standard youth-match depletion curves (not match-specific data)
Second-half observations were used only for retrospective comparison, not for generating recommendations.
Leakage Safeguards
  1. 1.
    No outcome-based tuning: The final score (4:3) was not used in any DSS computation or parameter selection.
  2. 2.
    No iterative refinement: The recommendation was generated in a single pass; no adjustments were made after observing the result.
  3. 3.
    Blind expert evaluation: Expert reviewers assessed the recommendation without knowledge of the match outcome.
6.1.6 Reproducibility Configuration
All computations for this pilot study are reproducible using the public repository:
python compute_pilot_distances.py
Key configuration parameters:
  • Random seed: SEED = 41 (same as synthetic experiments)
  • Categorical mapping: Hoch $\to 0.85$, Mittel $\to 0.50$, Niedrig $\to 0.20$
  • Fatigue projection: $A_{8}^{\text{proj}}=A_{8}^{\text{HT}}-0.15$
  • Missing attributes: Excluded from distance computation (reduced 5-dimensional space)
  • Opponent modeling: Disabled ($\alpha=0$) due to absence of opponent data
  • Weight configuration: Default dynamic weights as specified in Section 3.6.2.
@@ 6.2 Data Source and Match Context
  • Match duration: 2 $\times$ 35 min
@@ 6.3 Observed Attributes and Mapping Protocol
Match observers recorded six team attributes at the conclusion of each half using German terminology consistent with the source data collection protocol. To ensure full traceability between raw observations and DSS computations, we establish a systematic localization policy and provide a complete mapping specification.
6.3.1 Localization Policy
All DSS outputs, diagnostic reports, and experimental results presented in this paper use the canonical English attribute identifiers ($A_{1}$$A_{14}$) as defined in Table 2. When working with non-English source data:
  1. 1.
    Input normalization: Source attributes are mapped to the corresponding $A_{j}$ identifier using the translation table below. This mapping is applied at data ingestion, before any computation.
  2. 2.
    Internal representation: All internal computations use the English identifiers exclusively. The weight vector $w$, distance calculations, and diagnostic deltas reference $A_{1}$$A_{14}$.
  3. 3.
    Output standardization: All figures, tables, and textual outputs use English identifiers with German source terms provided parenthetically where relevant for auditability.
  4. 4.
    Categorical value translation: The German three-level scale (Hoch/Mittel/Niedrig) is converted to numerical values as specified in Equation (14).
6.3.2 German–English Attribute Mapping
Table 15 provides the complete one-to-one mapping between German source terms, English DSS identifiers, definitions, and computation methods for the pilot study data.
Table 15: Complete mapping of German observed attributes to DSS semantic space. All six observed attributes map to five unique DSS dimensions.
German Term DSS ID English Name Definition & Computation
Offensivkraft $A_{1}$ Offensive Strength Capacity to create and convert scoring opportunities. Direct correspondence; categorical value mapped via Equation (14).
Kompakte Defensive $A_{2}$ Defensive Strength Ability to maintain defensive shape and prevent attacks. Direct correspondence; categorical mapping.
Direkte vertikale Angriffe $A_{4}$ Transition Speed Capability for rapid vertical progression. Combined with Gegenangriff via $\max(\cdot)$ aggregation.
Gegenangriff $A_{4}$ Transition Speed Counterattacking capability after regaining possession. Combined with Direkte vertikale Angriffe.
Gegenpressing $A_{5}$ High Press Capability Aptitude for immediate pressure after ball loss. Direct correspondence; categorical mapping.
Restenergie $A_{8}$ Residual Energy Current stamina reserves. Direct correspondence; categorical mapping.
Aggregation Rule for $A_{4}$
Two German terms (Direkte vertikale Angriffe and Gegenangriff) both capture aspects of transition play. For DSS computation, these are aggregated to a single $A_{4}$ value:
$A_{4}=\max\bigl(v(\text{Direkte vertikale Angriffe}),\;v(\text{Gegenangriff})\bigr)$
where $v(\cdot)$ denotes the categorical-to-continuous conversion. This $\max$ aggregation reflects the tactical intuition that transition capability is demonstrated by either vertical directness or counterattacking effectiveness.
Unmapped Attributes
The pilot observation protocol captured 6 German attributes that map to 5 unique DSS dimensions ($A_{1}$, $A_{2}$, $A_{4}$, $A_{5}$, $A_{8}$). The remaining 9 attributes ($A_{3}$, $A_{6}$, $A_{7}$, $A_{9}$, $A_{10}$, $A_{11}$, $A_{12}$, $A_{13}$, $A_{14}$) were not directly observed and are therefore excluded from the reduced-dimension analysis in Section 6.5. Future validation studies should employ expanded observation protocols to achieve full attribute coverage.